# EnvGS: Modeling View-Dependent Appearance with Environment Gaussian

### v1

@@ Abstract
* Equal Contribution.
@@ 2 Related Work
Our work falls into the area of learning scene representations from a set of posed RGB images. In this section, we review both NeRF-based and Gaussian Splatting-based representations, with a focus on how these methods handle view-dependent effects.
@@
Recently, 3D Gaussian Splatting (3DGS) [15] has made significant strides toward real-time novel view synthesis. Unlike volume-based rendering, Gaussian Splatting uses efficient rasterization to render spatial Gaussian kernels, achieving real-time performance at 1080P resolution. Mip-Splatting [44] further introduces a 3D smoothing filter and a 2D Mip filter for alias-free renderings. Scaffold-GS [22] introduces structured 3D Gaussians to improve rendering efficiency and quality. [5] introduces triple splitting for efficient relighting. 2D Gaussian Splatting (2DGS)  [11] replaces the 3D Gaussian kernels with 2D Gaussian that better align with 3D surfaces, leading to more accurate surface reconstruction. However, these methods face challenges in accurately modeling reflections due to their use of spherical harmonics (SH) to parameterize view-dependent effects based solely on viewing direction, which often results in blurry reflections. More recent works GaussianShader [12] and 3DGS-DR [16] try to incorporate additional environment maps to improve reflection modeling ability. However, these methods only consider distant lighting, ignoring near-field lighting, and are unable to capture high-frequency reflection details. 3iGS [32] extends 3D Gaussian Splatting by augmenting it with an illumination field using tensorial factorization, rendering final reflections through a neural renderer. However, this method is limited to bounded scenes, which restricts its applicability in real-world, unbounded environments. Our method adopts 2DGS as the scene representation and models environmental illumination using an additional set of environment Gaussian, which is rendered via our proposed Gaussian tracer. Our approach inherently supports high-frequency, near-field, and distant lighting in unbounded scenes, enabling detailed reflection rendering while maintaining real-time performance.
@@ 3 Preliminary
$\mathbf{c}=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},$ (2)
with
$\alpha_{i}=\sigma_{i}\mathcal{G}_{i}\ \textrm{and}\ T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
@@
Compared to 3D Gaussian, 2D Gaussian offers distinct advantages as a surface representation. First, the ray-splat intersection method adopted by 2DGS avoids multi-view depth inconsistency. Second, 2D Gaussian inherently provides a well-defined normal, which is essential for high-quality surface reconstruction and accurate reflection calculations. However, 2DGS relies on the limited representational capacity of Spherical Harmonics (SH) to model scene appearance, preventing it from capturing strong view-dependent effects like specular reflections, which leads to poor rendering results and “foggy” geometry [33]. To this end, we use the geometry-aligned 2D Gaussian primitives as the base scene representation and demonstrate how we effectively model complex reflections in the next section.
@@ 5 Experiments
Methods Ref-NeRF Real Scenes NeRF-Casting Shiny Scenes FPS $\uparrow$
@@
Real-time 3DGS [15] 23.700 0.641 0.262 28.860 0.877 0.159 182.307
2DGS [11] 23.804 0.654 0.281 28.276 0.862 0.193 159.188
GaussianShader [12] 22.875 0.622 0.314 26.412 0.835 0.216 27.945
3DGS-DR [43] 23.522 0.640 0.274 28.487 0.858 0.197 133.593
Ours 24.617 0.671 0.241 30.444 0.886 0.148 26.221
Table 1: Quantitative comparison on Ref-Real [33] and NeRF-Casting Shiny Scenes [34] datasets.. Our method delivers the highest rendering quality among real-time techniques and outperforms several non-real-time methods, achieving competitive results with the state-of-the-art non-real-time method NeRF-Casting [34], while being 100 times faster. Note that Ref-NeRF* is an improved version of Ref-NeRF [33] that uses Zip-NeRF’s geometry model. All metrics are evaluated at 1/4 resolution as prior works  [34].
@@ 5.3 Baseline Comparisons
We compare our method with both implicit and explicit prior works, including Zip-NeRF [11], 3DGS [15], 2DGS [11], which are designed for general scenes with primarily diffuse appearances, as well as with methods specifically tailored for scenes with strong specular reflections, including Ref-NeRF [33], GaussianShader [12] and 3DGS-DR [16]. We also compare our approach with the more recent implicit method NeRF-Casting [34].
@@
Ref-NeRF [33] 27.078 0.761 0.341 23.620 0.673 0.389
3DGS [15] 26.406 0.835 0.168 25.300 0.882 0.165
2DGS [11] 26.057 0.831 0.190 24.933 0.876 0.184
GaussianShader [12] 25.113 0.809 0.216 23.061 0.840 0.226
3DGS-DR [43] 26.368 0.831 0.191 23.943 0.850 0.210
Ours 27.604 0.847 0.151 25.599 0.885 0.151
@@
We provide additional comparisons on the Mip-NeRF 360 dataset [3] and the Shiny Blender dataset [33] in the supplementary material, which also includes per-scene metric breakdowns from Tab. 1. Our method achieves competitive results on both datasets, demonstrating its versatility and effectiveness in handling diverse real-world scenes with complex view-dependent effects.
@@ 6 Conclusion and Discussion
In the supplementary material, we provide more qualitative and quantitative results and per-scene breakdowns to demonstrate the effectiveness and robustness of our method (Sec. A). We also provide additional ablation studies to further analyze the key components of our method (Sec. B). Furthermore, we provide details on the gradient computation of our Gaussian tracer (Sec. C).
@@ Appendix A Additional Results
A.1 Comparisons on Real-World Reflective Scenes
We present additional qualitative comparisons on the NeRF-Casting Shiny Scenes [34], which include both indoor and outdoor real-world scenes featuring complex reflections. As shown in Fig. 7, our method significantly outperforms previous approaches in reflection fidelity and overall rendering quality, particularly excelling in near-field reflections and high-frequency reflection details.
@@ A.1 Comparisons on Real-World Reflective Scenes
A.2 Comparisons on Shiny Blender [33]
@@ B.2 Speed Analysis
Rendering speed analysis. As mentioned in Sec. 4.1, the rendering of our method consists of two main rounds: rasterization of the base Gaussian and ray tracing of the environment Gaussian, the final color is the blending of the two. Based on the fact that only a small portion of the scene surface contains strong specular reflections, we can further accelerate the rendering process by only tracing rays with high blending weights, which are only made possible by our tracing-based renderer. We ablate the effectiveness and quality impact of this acceleration technique, results are shown in Tab. 4.
@@
Appendix C Details of Gradient Computation
@@
PSNR $\uparrow$ Ref-Real [33] NeRF-Casting Shiny Scenes [34]
@@
SSIM $\uparrow$ Ref-Real [33] NeRF-Casting Shiny Scenes [34]
@@
LPIPS $\downarrow$ Ref-Real [33] NeRF-Casting Shiny Scenes [34]
@@
PSNR $\uparrow$ Shiny Blender Scenes [33]
@@
SSIM $\uparrow$ Shiny Blender Scenes [33]
@@
LPIPS $\downarrow$ Shiny Blender Scenes [33]
@@
PSNR $\uparrow$ Mip-NeRF 360 [3]
@@
SSIM $\uparrow$ Mip-NeRF 360 [3]
@@
LPIPS $\downarrow$ Mip-NeRF 360 [3]

### v2

@@ Abstract
* Equal Contribution. $\dagger$ Corresponding author: Xiaowei Zhou
@@ 2 Related Work
In computer vision and graphics research [50, 9, 46, 12, 20, 40, 23], our work falls into the area of learning scene representations from a set of posed RGB images. In this section, we review NeRF-based and Gaussian Splatting-based methods, particularly their handling of view-dependent effects.
@@
Recently, 3D Gaussian Splatting (3DGS) [17] has made significant strides toward real-time novel view synthesis. Unlike volume-based rendering, Gaussian Splatting uses efficient rasterization to render spatial Gaussian kernels, achieving real-time performance at 1080P resolution. Mip-Splatting [52] further introduces a 3D smoothing filter and a 2D Mip filter for alias-free renderings. Scaffold-GS [26] introduces structured 3D Gaussians to improve rendering efficiency and quality. [5] introduces triple splitting for efficient relighting. 2D Gaussian Splatting (2DGS)  [13] replaces the 3D Gaussian kernels with 2D Gaussian that better align with 3D surfaces, leading to more accurate surface reconstruction. However, these methods face challenges in accurately modeling reflections due to their use of spherical harmonics (SH) to parameterize view-dependent effects based solely on viewing direction, which often results in blurry reflections. More recent works GaussianShader [14] and 3DGS-DR [18] try to incorporate additional environment maps to improve reflection modeling ability. However, these methods only consider distant lighting, ignoring near-field lighting, and are unable to capture high-frequency reflection details. 3iGS [36] extends 3DGS by augmenting it with an illumination field using tensorial factorization, rendering final reflections through a neural renderer. However, it is limited to bounded scenes, which restricts its applicability in real-world, unbounded environments. Concurrent work [54] choose to model near-field reflection with tensorial factorization and far-field with spherical feature grid. Our method adopts 2DGS as the scene representation and models environmental illumination using an additional set of environment Gaussian, which is rendered via our proposed Gaussian tracer. Our approach inherently supports high-frequency, near-field, and distant lighting in unbounded scenes, enabling detailed reflection rendering while maintaining real-time performance.
@@ 3 Preliminary
$\mathbf{c}=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},\ \textrm{with}\ \alpha_{i}=\sigma_{i}\mathcal{G}_{i},\ T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
@@
Compared to 3D Gaussian, 2D Gaussian offers distinct advantages as a surface representation. First, the ray-splat intersection method adopted by 2DGS avoids multi-view depth inconsistency. Second, 2D Gaussian inherently provides a well-defined normal, which is essential for high-quality surface reconstruction and accurate reflection calculations. However, 2DGS relies on the limited representational capacity of Spherical Harmonics (SH) to model scene appearance, preventing it from capturing strong view-dependent effects like specular reflections, which leads to poor rendering results and “foggy" geometry [37]. To this end, we use the geometry-aligned 2D Gaussian primitives as the base scene representation and demonstrate how we effectively model complex reflections in the next section.
@@ 5 Experiments
Methods Ref-NeRF Real Scenes [37] NeRF-Casting Scenes [38] FPS $\uparrow$ Training Time $\downarrow$
@@
Real-time 3DGS [17] 23.700 0.641 0.262 28.860 0.877 0.159 182.307 0.6h
2DGS [13] 23.804 0.654 0.281 28.276 0.862 0.193 159.188 0.7h
GaussianShader [14] 22.875 0.622 0.314 26.412 0.835 0.216 27.945 1.6h
3DGS-DR [51] 23.522 0.640 0.274 28.487 0.858 0.197 133.593 1.0h
Ours 24.617 0.671 0.241 30.444 0.886 0.148 26.221 2.5h
Table 1: Quantitative comparison on Ref-Real [37] and NeRF-Casting Scenes [38] datasets. Our method delivers the highest rendering quality among real-time techniques and outperforms several non-real-time methods, achieving competitive results with the state-of-the-art non-real-time method NeRF-Casting [38], while being 100 times faster. Note that Ref-NeRF* is an improved version of Ref-NeRF [37] that uses Zip-NeRF’s geometry model. All metrics are evaluated at 1/4 resolution as prior works  [38].
@@ 5.1 Implementation Details
Methods Foreground Near-Field
[…]
3DGS [17] 31.681 0.955 0.046 44.161 0.994 0.008
2DGS [13] 31.328 0.953 0.053 43.624 0.994 0.010
GaussianShader [14] 30.694 0.947 0.058 42.391 0.993 0.010
3DGS-DR 31.814 0.953 0.050 44.007 0.994 0.009
Ours 33.295 0.962 0.040 46.392 0.996 0.007
Table 2: Quantitative results of foreground and near-field region on Ref-Real [37] and NeRF-Casting Scenes [38]. See supplementary Sec. A.1 for more details and qualitative results.
@@ 5.3 Baseline Comparisons
We compare our method with both implicit and explicit prior works, including Zip-NeRF [13], 3DGS [17], 2DGS [13], which are designed for general scenes with primarily diffuse appearances, as well as with methods specifically tailored for scenes with strong specular reflections, including Ref-NeRF [37], GaussianShader [14], 3DGS-DR [18], and NeRF-Casting [38]. We also compare with ENVIDR [22], NDE [45] these two object-level methods.
@@
Ref-NeRF [37] 24.529 0.713 0.397 23.620 0.673 0.389
3DGS [17] 26.171 0.825 0.181 25.300 0.882 0.165
2DGS [13] 25.869 0.819 0.198 24.933 0.876 0.184
GaussianShader [14] 24.768 0.795 0.225 23.061 0.840 0.226
3DGS-DR [51] 26.093 0.820 0.192 23.943 0.850 0.210
Ours 28.629 0.876 0.121 25.599 0.885 0.151
@@
We provide additional comparisons on the Mip-NeRF 360 dataset [3], the Shiny Blender dataset [37] and with ENVIDR [22], NDE [45] in supplementary Appendix A, which also includes per-scene metric breakdowns from Tab. 1. Our method achieves competitive results on both datasets, demonstrating its versatility and effectiveness in handling diverse scenes with complex view-dependent effects.
@@ 5.4 Ablation Studies
Perceptual loss. The “w/o lpips loss” variants removes the perceptual loss. The results in Tab. 4 demonstrate that our method still produces detailed and accurate reflections without perceptual loss, which offers only marginal improvements.
@@ 6 Conclusion and Discussion
Acknowledgements. This work was partially supported by NSFC (No. U24B20154, 62402427), Ant Group, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@
In the supplementary material, we provide more qualitative and quantitative results and per-scene breakdowns to demonstrate the effectiveness and robustness of our method (Sec. A). We also provide additional ablation studies to further analyze the key components of our method (Sec. B). Furthermore, we provide details on the gradient computation of our Gaussian tracer (Sec. D).
@@ Appendix A Additional Results
A.1 Comparison on Reflective Regions
To demonstrate the improvements in the reflective and near-field reflection regions using our environment Gaussian representation, we additionally annotate a reflection mask to compute metrics specifically for the reflective region and a near-field mask to evaluate near-field reflections on the Ref-Real [37] and NeRF-Casting [38] datasets. As shown in Tab. 2 and Fig. 6, our method achieves a significant improvement of over 1.0 PSNR $\uparrow$ improvement on the reflective region and 2.0 PSNR $\uparrow$ in the near-field regions compared to using an environment map. These results highlight the effectiveness of our approach in capturing and rendering complex reflective and near-field phenomena.
Figure 6: Qualitative comparison on reflective foreground and near-field reflection regions. We also provide visualizations of the foreground and near-field region mask we annotated.
The reflective masks mentioned above are obtained through the following steps. First, we train our EnvGS on each scene, then export the trained Gaussian and remove the Gaussian points in 3D space except for those in the foreground reflective region. We render the remaining Gaussian to generate an accumulated alpha map. Finally, we binarize this alpha map to obtain the foreground reflective masks. We manually annotate the near-field masks as they are difficult to define in 3D space.
The quantitative results in Tab. 2 are evaluated only in the masked regions, following NeRF-Casting [38], we compute these masked metrics by blending the masked regions onto a white background.
A.2 Comparison on Real-World Shiny Scenes
We present additional qualitative comparisons on the NeRF-Casting Shiny Scenes [38], including both indoor and outdoor real-world scenes featuring complex reflections. As shown in Fig. 9, our method significantly outperforms previous approaches in reflection fidelity and overall rendering quality, particularly excelling in near-field reflections and high-frequency reflection details.
@@ A.2 Comparison on Real-World Shiny Scenes
A.3 Comparison on Shiny Blender [37]
@@ A.4 Comparison on Mip-NeRF 360 [3]
Methods Ref-Real [37] and NeRF-Casting [38]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
ENVIDR 15.890 0.416 0.607 0.058
NDE 19.399 0.422 0.593 0.083
Ours 27.947 0.794 0.189 26.221
Table 5: Quantitative comparison with object-level methods.
Figure 7: Qualitative comparison with object-level methods.
A.5 Additional Baselines
We also compare our method with object-baselines including ENVIDR [22] and NDE [45]. While object-level methods perform well on synthetic data, they often struggle with real-world scenes and cannot real-time rendering speed on scenes with background, as shown in Tab. 5 and Fig. 7.
@@ B.2 Speed Analysis
Rendering speed analysis. As mentioned in Sec. 4.1, the rendering of our method consists of two main rounds: rasterization of the base Gaussian and ray tracing of the environment Gaussian, and the final color is the blending of the two. Based on the fact that only a small portion of the scene surface contains strong specular reflections, we can further accelerate the rendering process by only tracing rays with high blending weights, which are only made possible by our tracing-based renderer. We ablate the effectiveness and quality impact of this acceleration technique, results are shown in Tab. 6.
B.3 Environment Gaussian Design
@@
Appendix C Details of Environment Gaussian
We provide more details of our base Gaussian and environment Gaussian. The SH coefficients of both base Gaussian and environment Gaussian are set to two for the best results. The environment Gaussian is jointly optimized with the base Gaussian, and environment Gaussian constitutes around 15% of the base Gaussian, of average 300k Gaussian primitives using 70MB after training. For pruning, we follow the pruning method in the original 2DGS [13] and keep at most the top 630k environment Gaussian primitives based on rendering weights.
Appendix D Details of Gradient Computation
@@
PSNR $\uparrow$ Methods Ref-Real [37] NeRF-Casting Shiny Scenes [38]
@@
SSIM $\uparrow$ Methods Ref-Real [37] NeRF-Casting Shiny Scenes [38]
@@
LPIPS $\downarrow$ Methods Ref-Real [37] NeRF-Casting Shiny Scenes [38]
@@
PSNR $\uparrow$ Methods Shiny Blender Scenes [37]
@@
SSIM $\uparrow$ Methods Shiny Blender Scenes [37]
@@
LPIPS $\downarrow$ Methods Shiny Blender Scenes [37]
@@
PSNR $\uparrow$ Methods Mip-NeRF 360 [3]
@@
SSIM $\uparrow$ Methods Mip-NeRF 360 [3]
@@
LPIPS $\downarrow$ Methods Mip-NeRF 360 [3]